OpenStudy (anonymous):

Solve the following system of equations. 2x + 2y + z = 10 3x – y + 3z = 10 2x + 3y – 2z = 6 (2, 2, 2) (–2, 2, 2) (2, –2, 2) (2, 2, –2)

OpenStudy (anonymous):

im not sure if im doing this right im about to put up my work in a sec

OpenStudy (abbie):

can you post this in the mathematics section?

OpenStudy (anonymous):

lol omg i didnt even notice that

OpenStudy (abbie):

look for examples

OpenStudy (abbie):

NP

OpenStudy (abbie):

heres an example I found (1) 2x + 3y - z = 4 (2) 3x + 2y + 2z = 6 (3) 4x - y + 3z = 5 multiply (3) by 2 8x - 2y + 6z = 10 (2) 3x + 2y + 2z = 6 add them 11x + 8z = 16 multiply (3) by 3 (3) 12x - 3y + 9z = 15 (1) 2x + 3y - z = 4 add them 14x + 8z = 19 11x + 8z = 16 subtract 14x + 8z = 19 -11x -8z = -16 3x = 3 x = 1 ------------- 11x + 8z = 16 11*1 +8z = 16 8z = 5 z = 5/8 --------------